OpenStudy (anonymous):

Prove that the given rule is a linear transformation, or show why it isn't. T : C^∞(R)→R, T(f)= b∫a (f)(sin(x))dx a = 1(bottom), b = 3 (top)

OpenStudy (anonymous):

\[T(f)=\int_a^bf(x)\sin(x)dx\] need to check that \[T(f+g)=T(f)+T(g)\] i.e. \[\int_a^b(f(x)+g(x))\sin(x)dx=\int_a^bf(x)\sin(x)dx+\int_a^bg(x)\sin(x)dx\] which i believe it is

OpenStudy (anonymous):

also that \[T(cf)=cT(f)\] i.e. \[\int_a^bcf(x)\sin(x)dx=c\int_a^bf(x)\sin(x)dx\] again pretty clearly yes

OpenStudy (anonymous):

oh are your functions integrable?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

ahah, so wait ... none of that applies then?